Use the References to access important values if needed for this question. Given the standard enthalpy changes for the follow(please bold or circle the answer that I need to input)
Answer #1

Lets number the reaction as 1, 2, 3 from top to bottom
required reaction should be written in terms of other reaction
This is Hess Law
required reaction can be written as:
reaction 3 = -0.5 * (reaction 1) +0.5 * (reaction 2)

So, ΔHo rxn for required reaction will be:
ΔHo rxn = -0.5 * ΔHo rxn(reaction 1) +0.5 * ΔHo rxn(reaction 2)
= -0.5 * (-434.6) +0.5 * (-696.6)
= -131.0 KJ
Answer: -131.0 KJ